OpenStudy (hrhsbull):

Indicate the equation of the line through (2, -4) and having slope of 3/5 (im desperate)

OpenStudy (anonymous):

Use this : ;) Y2-Y=m(X2-X) In there we have : Y2=-4 X2=2

OpenStudy (anonymous):

@HrhsBull :Understoond ???:)

OpenStudy (hrhsbull):

thanks